# Orthogonal roots, quantum Hafnians, and generalized Rothe diagrams

### v1

@@ 2.3. Coplanar quadruples and $D_{4}$-subsystems
The following proposition expands [8, Proposition 3.3] and details many useful properties of coplanar quadruples.
@@
Part (i) follows from [8, Proposition 3.8] and the first paragraph of its proof.
@@
Definition 2.12 resembles but differs from [8, Definition 3.9], where the classification of coplanar quadruples into the three types of possible features is characterized in a less concise way. However, the definitions are equivalent: in the notation of Example 2.7, each coplanar quadruple is necessarily one of $Q_{1},Q_{2}$ and $Q_{3}$, and both definitions classify $Q_{1}$ as an alignment, $Q_{2}$ as a crossing, and $Q_{3}$ as an alignment. In other words, both definitions guarantee that each coplanar quadruple $Q$ is an alignment, crossing, or nesting if and only if the perfect matching corresponding to it is a feature of the same type under the usual coordinates for the $D_{4}$-subsystem $\Psi=\text{\rm Span}(Q)\cap\Phi$ associated to $Q$.
@@
Part (i) and (ii) a restatement of [8, Proposition 3.20]. Part (iii) follows from [8, Lemma 3.18], which states that the features of $R$ form the blocks of a Steiner quadruple system $S(3,4,8)$, meaning that every $3$-element subset of $R$ is contained in a unique feature. More precisely, there is a unique Steiner quadruple system on $8$ points ([8, Remark 3.16]) and its blocks are closed under taking complements, so (iii) follows. ∎
3. Residues and quasiparabolic sets
In Section 3.1, we establish the key new definitions of residues, Rothe diagrams, the level statistic, as well as generalized Hafnians and Pfaffians and their $q$-analogues. We also recall the notion of a quasiparabolic set, due to Rains and Vazirani. Then, in Section 3.2, we give details of two of the motivating examples of residues: one based on perfect matchings of a set of size $2k$ (Theorem 3.10), and one based on permutations of $k$ objects (Theorem 3.13).
@@ 3.1. Key definitions
The following definition introduces the key notion of residues, which applies to all root systems. This notion is new to the best of our knowledge.
@@
    For any orthogonal set $R\in\Omega_{U}$, we define the set of residues of $R$ with respect to $U$ to be
    $\text{\rm Res}_{U}(R)=\{\gamma\in U:s_{\beta}(\gamma)>0\text{\ for\ all\ }\beta\in R\},$
    and we will also refer to $\text{\rm Res}_{U}(R)$ as the (generalized) Rothe diagram of $R$ with respect to $U$. We define $\rho_{U}(R):=\lvert\text{\rm Res}_{U}(R)\rvert$ to be the level of $R$ with respect to $U$.
@@
    We identify an element $R=\{\beta_{1},\beta_{2},\ldots,\beta_{k}\}\in\Omega_{U}$ with the monomial $\prod_{\beta\in R}x_{\beta}$, where $\{x_{\beta}:\beta\in U\}$ is a set of commuting indeterminates. The (generalized) quantum Hafnian of $U$, $\text{\rm QHf}(U)=\text{\rm QHf}(U,q)$, is the element of ${\mathbb{Z}}[U][q]$ given by
@@
    The specializations $\text{\rm QPf}(U):=\text{\rm QHf}(U,-q)$, $\text{\rm Hf}(U):=\text{\rm QHf}(U,1)$, and $\text{\rm Pf}(U):=\text{\rm QHf}(U,-1)$ are called the (generalized) quantum Pfaffian, the (generalized) Hafnian, and the (generalized) Pfaffian of $U$, respectively.
@@
    Since $s_{\gamma}(\gamma)=-\gamma<0$ for any positive root $\gamma$ in a root system, it follows that $\text{\rm Res}_{U}(R)$ cannot contain any element of $R$ in the setting of Definition 3.1.
@@
$\text{\rm Res}_{U}(Q_{1})=\emptyset,\quad\text{\rm Res}_{U}(Q_{2})=\{\theta\}=\{\varepsilon_{1}+\varepsilon_{2}\},\quad\text{and}\quad\text{\rm Res}_{U}(Q_{3})=\{\theta-\alpha_{2},\ \theta\}=\{\varepsilon_{1}+\varepsilon_{2},\ \varepsilon_{1}+\varepsilon_{3}\},$
@@
Note that the sets $Q_{i}$ for $1\leq i\leq 3$ are exactly the three coplanar quadruples partitioning $\Phi_{+}$} from Example 2.7, with $Q_{1}$ being an alignment, $Q_{2}$ a crossing, and $Q_{3}$ a nesting. The highest root $\theta$ and the second highest root $\theta-\alpha_{2}$ are the only roots that may appear in $\text{\rm Res}_{U}(R)$ for any $R\in\Omega_{U}$. Note also that if $\alpha$ is a residue and $\alpha\in\{\theta,\theta-\alpha_{2}\}$, then $\alpha$ has full support with respect to any of the orthogonal bases $Q_{i}=\{\beta_{1},\beta_{2},\beta_{3},\beta_{4}\}$.
The next definition considers an important special case of Definition 3.1 (ii). Note that by Section 2.1, if $R$ is a $k$-root of a root system $\Phi$, then the set $\Psi:=\text{\rm Span}(R)\cap\Phi$ is a subsystem of $\Phi$, and the positive roots $\Phi_{+}$ of $\Phi$ induce a positive system $\Psi_{+}=\Psi\cap\Phi_{+}=\text{\rm Span}(R)\cap\Phi_{+}$ of $\Psi$.
@@
Let $\Phi$ be an arbitrary root system, let $R\subseteq\Phi_{+}$ be a $k$-root of $\Phi$, and let $\Psi_{R}=\text{\rm Span}(R)\cap\Phi$ be the subsystem of $\Phi$ induced by $R$. We define the set of residues of $R$ to be the set
$\text{\rm Res}(R)=\text{\rm Res}_{\Psi_{+}}(R),$
where $\Psi_{+}=\text{\rm Span}(R)\cap\Phi_{+}$ is the positive system of $\Psi$ induced by $\Phi_{+}$. We refer to $\text{\rm Res}(R)$ as the (generalized) Rothe diagram of $R$, and we define $\rho(R):=\lvert\text{\rm Res}(R)\rvert$ to be the level of $R$.
@@
    It follows from Definitions 3.1 (ii) and 3.4 that for any $k$-root $R$ of a root system $\Phi$, we have $\text{\rm Res}(R)=\text{\rm Res}_{\Phi_{+}}(R)\cap\text{\rm Span}(R)$. In particular, if $R$ is an $n$-root where $n=\text{\rm rank}(\Phi)$, then $\text{\rm Res}(R)=\text{\rm Res}_{\Phi_{+}}(R)$ and $\rho(R)=\rho_{\Phi_{+}}(R)$.
@@
    In the setting of Example 3.3, we have $\text{\rm Res}(Q_{i})=\text{\rm Res}_{\Phi_{+}}(Q_{i})=\text{\rm Res}_{U}(Q_{i})$ and $\rho(Q_{i})=\rho_{\Phi_{+}}(Q_{i})=\rho_{U}(Q_{i})$ for each $i$. Since the calculation of the residue sets in this example depends only on the choice of a simple system for a type $D_{4}$ root system, it follows that whenever $Q$ is a feature in a root system of type $E_{7},E_{8}$ or $D_{n}$ for $n$ even, we have $\rho(Q)=0$ if $Q$ is an alignment, $\rho(Q)=1$ if $Q$ is a crossing, and $\rho(Q)=2$ if $Q$ is a nesting.
Quasiparabolic sets were introduced by Rains and Vazirani for a general Coxeter system as follows.
@@
[14, Section 2, Section 5] Let $(W,S)$ be a Coxeter system with set of reflections $T$. A scaled $W$-set is a pair $(X,\lambda)$ where $X$ is a $W$-set and $\lambda:X\rightarrow{\mathbb{Z}}$ is a function satisfying $|\lambda(sx)-\lambda(x)|\leq 1$ for all $s\in S$. An element $x\in X$ is $W$-minimal if $\lambda(sx)\geq\lambda(x)$ for all $s\in S$ and is $W$-maximal if $\lambda(sx)\leq\lambda(x)$ for all $s\in S$.
@@
For any quasiparabolic $W$-set $(X,\lambda)$, it follows from [14, Corollary 2.10] that each $W$-orbit in $X$ contains at most one $W$-minimal element, so if $X$ is a finite and transitive $W$-set then $X$ has a unique $W$-minimal element, $x_{0}$. Moreover, when this is the case, the level of any element $x\in X$ equals the minimum length of the elements $w\in W$ such that $w(x_{0})=x$ by [14, Corollary 2.13].
We may associate a quantum Hafnian and a Poincaré polynomial to a general quasiparabolic set by defining $\text{\rm QHf}(X)=\sum_{x\in X}q^{\rho(R)}x$ and $PS_{X}(q)=\sum_{x\in X}q^{\rho(x)}.$ However, we will not require this level of generality, because our focus for the rest of the paper will be on quasiparabolic sets that arise from sets of orthogonal roots via the following construction.
@@ 3.2. Motivating examples
We now present the two motivating examples of residues: one based on perfect matchings, and the other based on permutations. The following definition will be helpful in the context of the perfect matching example.
@@
Let $M=\{\{i_{1},j_{1}\},\ldots,\{i_{k},j_{k}\}\}$ be a perfect matching of the set $[2k]$, where we have $k\geq 2$ and $i_{1}<\ldots<i_{k}$ and $i_{t}<j_{t}$ for all $1\leq t\leq k$. Following [12, §2], we define $\sigma_{M}\in S_{2k}$ to be the permutation whose two-line notation is given by
@@
We define $\tau_{M}\in S_{2k}$ to be the fixed point free involution whose disjoint cycle notation is given by
@@
    $\{(i,j):j<i\text{\ and\ }u_{ji}\in\text{\rm Res}_{U}(R(M))\}=\{(i,j):(j,i)\in D_{M}\}$
    is the Rothe diagram $\hat{D}_{\tt FPF}(\tau_{M})$ of the fixed point free involution $\tau_{M}$ in the sense of Hamaker–Marberg–Pawlowski [9, (3.3)].
@@
    The generalized quantum Hafnian $\text{\rm QHf}(U)$ is the $q$-Hafnian in the sense of Jing–Zhang [13, Proposition 3.6], and the generalized quantum Pfaffian $\text{\rm QPf}(U)$ is the $q$-Pfaffian in the sense of Strickland [16] and Jing–Zhang [13, Proposition 3.4].
@@
    The Hafnian $\text{\rm Hf}(U)$ is the Hafnian of a $2k\times 2k$ symmetric matrix $(a_{ij})$ satisfying $a_{ij}=u_{ij}$ for all $1\leq i<j\leq 2k$, and the Pfaffian $\text{\rm Pf}(U)$ is the Pfaffian of a $2k\times 2k$ skew-symmetric matrix $(a_{ij})$ satisfying $a_{ij}=u_{ij}$ for all $1\leq i<j\leq 2k$.
@@
If $\beta,u_{ij}\in U$ satisfy $s_{\beta}(u_{ij})<0$, then we must either have $\beta=u_{aj}$ for some $a\leq i$, or $\beta=u_{ib}$ for some $b\leq j$. It follows that $u_{ij}$ is a residue of $R(M)$ if and only if we have both $\tau_{M}(i)>j$ and $\tau_{M}(j)>i$, proving (iii).
@@
is the Rothe diagram $\hat{D}_{\tt FPF}(\tau)$ of the fixed point free involution $\tau$ as defined by Hamaker–Marberg–Pawlowski [9, (3.3)]. Part (iv) follows by applying (iii) to this definition.
It follows from [9, Proposition 3.6] that the level, $\rho_{U}(R(M))=|\hat{D}_{\tt FPF}(\tau_{M})|$, of $R$ is equal to $(\ell(\tau_{M})-k)/2$. Lemma 3.9 (iii) implies that $\rho_{U}(R(M))=\ell(\sigma_{M}$), and this completes the proof of (v).
Rains–Vazirani [14, §4] proved that the function $(\ell(\tau_{M})-k)/2$ is a quasiparabolic level function for the set of fixed point free involutions of $S_{2k}$, regarded as an $S_{2k}$-set in the natural way. Combining this result with part (v) proves (i).
@@
where the sum is over all perfect matchings of $[2k]$. This agrees with Jing–Zhang’s construction of the quantum Hafnian from [13, Proposition 3.6] All the assertions of (vi) and (vii) now follow from the relevant definitions.
Part (viii) is a standard result that has been rediscovered several times; see [8, Proposition 7.1]. ∎
@@
In certain cases, the product of the residues of a perfect matching, when regarded as a polynomial in the $\varepsilon_{i}$, has an interpretation in terms of Schubert polynomials; see [9, Theorem 1.3].
@@
    The generalized quantum Hafnian $\text{\rm QHf}(U)$ is the $q$-permanent of the matrix $A=(v_{ij})$ in the sense of [12, (2.7)], and the generalized quantum Pfaffian $\text{\rm QHf}(U)$ is the quantum row permanent $\text{\rm per}_{q}(A)$ of $A$ in the sense of [13, (3.18)].
@@
    The Hafnian $\text{\rm Hf}(U)$ and the Pfaffian $\text{\rm Pf}(U)$ are the permanent $\text{\rm per}(A)$, and the determinant $\det(A)$, respectively.
@@
Parts (ii) and (iii) follow by restriction from Theorem 3.10 (ii) and (iii), with each $k$-root $R(\pi)$ defined in (ii) corresponding to the perfect matching $M\{\{i,\pi(i)+k\}:i\in[k]\}$. Part (iv) then follows from the definition of the Rothe diagram $D(\pi)$ [9, (2.7)]. Part (v), which is well known, follows from the fact that there is a bijection between the inversions of $\pi$ and the Rothe diagram of $\pi$ given by $(i,j)\mapsto(i,\pi(j))$.
@@
By [14, Theorem 3.1], this makes $S_{k}$ into a quasiparabolic set with height function $\ell$, proving (i).
@@
    It is sometimes helpful to identify $U$ as the positions on a $k\times k$ board and $\Omega_{U}$ with configurations of $k$ nonattacking rooks on the board, with each root $v_{ij}=\varepsilon_{i}\pm\varepsilon_{j+k}$ corresponding to the $(i,j)$-th position. A position on the board then corresponds to a residue if and only if it is vacant and not attacked either from the left or from above. Plotting a $k$-root and its residues on the board yields the original Rothe diagram of the permutation correspoding to the rook configuration; see Figure 1.
@@
    Instead of regarding $\Omega_{U}$ as an $S_{k}\times S_{k}$-set, we may regard it as an $(S_{k},S_{k})$-biset, as in the proof of part (v) of the theorem, and from this perspective, $\Omega_{U}$ has a left $S_{k}$-action permuting the rows, and a right $S_{k}$-action permuting the columns. The actions commute with each other, and $\Omega_{U}$ is a quasiparaboic $S_{k}$-set under each action.

### v2

@@ 2.3. Coplanar quadruples and $D_{4}$-subsystems
The following proposition expands [12, Proposition 3.3] and details many useful properties of coplanar quadruples.
@@
Part (i) follows from [12, Proposition 3.8] and the first paragraph of its proof.
@@
Definition 2.13 resembles but differs from [12, Definition 3.9], where the classification of coplanar quadruples into the three types of possible features is characterized in a less concise way. However, the definitions are equivalent: in the notation of Example 2.8, each coplanar quadruple is necessarily one of $Q_{1},Q_{2}$ and $Q_{3}$, and both definitions classify $Q_{1}$ as an alignment, $Q_{2}$ as a crossing, and $Q_{3}$ as an alignment. In other words, both definitions guarantee that each coplanar quadruple $Q$ is an alignment, crossing, or nesting if and only if the perfect matching corresponding to it is a feature of the same type under the usual coordinates for the $D_{4}$-subsystem $\Psi=\text{\rm Span}(Q)\cap\Phi$ associated to $Q$.
@@
Parts (i) and (ii) are a restatement of [12, Proposition 3.20]. Part (iii) follows from [12, Lemma 3.18], which states that the features of $R$ form the blocks of a Steiner quadruple system $S(3,4,8)$, meaning that every $3$-element subset of $R$ is contained in a unique feature. More precisely, there is a unique Steiner quadruple system on $8$ points ([12, Remark 3.16]) and its blocks are closed under taking complements, so (iii) follows. ∎
3. Generalized Rothe diagrams and quasiparabolic sets
In Section 3.1, we introduce the definitions of generalized Rothe diagrams and several related objects, including generalized quantum Hafnians (Definition 3.2). We also recall Rains and Vazirani’s notion of quasiparabolic sets, which provides a common framework for the sets of $k$-roots that we will study in the remaining parts of the paper. In Section 3.2, we give details of two of the motivating examples of generalized Rothe diagrams: one based on perfect matchings of a set of size $2k$ (Theorem 3.11), and one based on permutations of $k$ objects (Theorem 3.14).
@@ 3.1. Key definitions
The following two definitions are central to this paper. They make sense for all root systems, and they are new to the best of our knowledge.
Definition 3.1.
Let $\Phi$ be an arbitrary root system and let $\alpha,\beta\in\Phi_{+}$. We say that $\alpha$ dominates $\beta$ if $s_{\beta}(\alpha)$ is a positive root.
@@
    For any orthogonal set $R\in\Omega_{U}$, we say that a positive root $\alpha\in\Phi_{+}$ dominates $R$, or that $\alpha$ is $R$-dominating, if $\alpha$ dominates every element of $R$. We define the generalized Rothe diagram of $R$ with respect to $U$ to be the set
    $\mathscr{D}_{U}(R)=\{\alpha\in U:\alpha\text{\;dominates\;}R\}=\{\alpha\in U:s_{\beta}(\alpha)>0\text{\ for\ all\ }\beta\in R\}.$
    We define the level of $R$ with respect to $U$ to be $\rho_{U}(R):=\lvert\mathscr{D}_{U}(R)\rvert$.
@@
    We associate to each element $R=\{\beta_{1},\beta_{2},\ldots,\beta_{k}\}\in\Omega_{U}$ the monomial $\prod_{\beta\in R}x_{\beta}$, where $\{x_{\beta}:\beta\in U\}$ is a set of commuting indeterminates. The generalized quantum Hafnian of $U$, $\text{\rm QHf}(U)=\text{\rm QHf}(U,q)$, is the element of ${\mathbb{Z}}[U][q]$ given by
@@
    The specializations $\text{\rm QPf}(U):=\text{\rm QHf}(U,-q)$, $\text{\rm Hf}(U):=\text{\rm QHf}(U,1)$, and $\text{\rm Pf}(U):=\text{\rm QHf}(U,-1)$ are called the generalized quantum Pfaffian, the generalized Hafnian, and the generalized Pfaffian of $U$, respectively.
@@
    Since $s_{\gamma}(\gamma)=-\gamma<0$ for any root $\alpha\in\Phi$, no positive root dominates itself, so the set $\mathscr{D}_{U}(R)$ cannot contain any element of $R$ in the setting of Definition 3.2.
@@
$\mathscr{D}_{U}(Q_{1})=\emptyset,\quad\mathscr{D}_{U}(Q_{2})=\{\theta\}=\{\varepsilon_{1}+\varepsilon_{2}\},\quad\text{and}\quad\mathscr{D}_{U}(Q_{3})=\{\theta-\alpha_{2},\ \theta\}=\{\varepsilon_{1}+\varepsilon_{2},\ \varepsilon_{1}+\varepsilon_{3}\},$
@@
Note that the sets $Q_{i}$ for $1\leq i\leq 3$ are exactly the three coplanar quadruples partitioning $\Phi_{+}$} from Example 2.8, with $Q_{1}$ being an alignment, $Q_{2}$ a crossing, and $Q_{3}$ a nesting. The highest root $\theta$ and the second highest root $\theta-\alpha_{2}$ are the only roots that may appear in $\mathscr{D}_{U}(R)$ for any $R\in\Omega_{U}$. Note also that if a root $\alpha\in\{\theta,\theta-\alpha_{2}\}$ dominates $Q_{i}$ for some $i$, then $\alpha$ has full support with respect to the orthogonal basis $Q_{i}$.
The next definition considers an important special case of Definition 3.2 (ii), where the reference set $U$ in Definition 3.2 no longer needs to be specified independently. Note that by Section 2.1, if $R$ is a $k$-root of a root system $\Phi$, then the set $\Psi:=\text{\rm Span}(R)\cap\Phi$ is a subsystem of $\Phi$, and the positive roots $\Phi_{+}$ of $\Phi$ induce a positive system $\Psi_{+}=\Psi\cap\Phi_{+}=\text{\rm Span}(R)\cap\Phi_{+}$ of $\Psi$.
@@
Let $\Phi$ be an arbitrary root system, let $R\subseteq\Phi_{+}$ be a $k$-root of $\Phi$, and let $\Psi_{R}=\text{\rm Span}(R)\cap\Phi$ be the subsystem of $\Phi$ induced by $R$. We defined the generalized Rothe diagram of $R$ to be the set
$\mathscr{D}(R)=\mathscr{D}_{\Psi_{+}}(R),$
where $\Psi_{+}=\text{\rm Span}(R)\cap\Phi_{+}$ is the positive system of $\Psi$ induced by $\Phi_{+}$. We define $\rho(R):=\lvert\mathscr{D}(R)\rvert$ to be the level of $R$.
@@
    It follows from Definitions 3.2 (ii) and 3.5 that for any $k$-root $R$ of a root system $\Phi$, we have $\mathscr{D}(R)=\mathscr{D}_{\Phi_{+}}(R)\cap\text{\rm Span}(R)$. In particular, if $R$ is an $n$-root where $n=\text{\rm rank}(\Phi)$, then $\mathscr{D}(R)=\mathscr{D}_{\Phi_{+}}(R)$ and $\rho(R)=\rho_{\Phi_{+}}(R)$.
@@
    In the setting of Example 3.4, we have $\mathscr{D}(Q_{i})=\mathscr{D}_{\Phi_{+}}(Q_{i})=\mathscr{D}_{U}(Q_{i})$ and $\rho(Q_{i})=\rho_{\Phi_{+}}(Q_{i})=\rho_{U}(Q_{i})$ for each $i$. Since the calculation of $\mathscr{D}(Q_{i})$ in this example depends only on the choice of a simple system for a type $D_{4}$ root system, it follows that whenever $Q$ is a feature in a root system of type $E_{7},E_{8}$ or $D_{n}$ for $n$ even, we have $\rho(Q)=0$ if $Q$ is an alignment, $\rho(Q)=1$ if $Q$ is a crossing, and $\rho(Q)=2$ if $Q$ is a nesting.
The rest of the paper will study sets of the form $\Omega_{U}$ that are quasiparabolic sets under their associated level functions, in the following sense.
@@
[19, Section 2, Section 5] Let $(W,S)$ be a Coxeter system with set of reflections $T$. A scaled $W$-set is a pair $(X,\lambda)$ where $X$ is a $W$-set and $\lambda:X\rightarrow{\mathbb{Z}}$ is a function satisfying $|\lambda(sx)-\lambda(x)|\leq 1$ for all $s\in S$. An element $x\in X$ is $W$-minimal if $\lambda(sx)\geq\lambda(x)$ for all $s\in S$ and is $W$-maximal if $\lambda(sx)\leq\lambda(x)$ for all $s\in S$.
@@
For any quasiparabolic $W$-set $(X,\lambda)$, it follows from [19, Corollary 2.10] that each $W$-orbit in $X$ contains at most one $W$-minimal element, so if $X$ is a finite and transitive $W$-set then $X$ has a unique $W$-minimal element, $x_{0}$. Moreover, when this is the case, the level of any element $x\in X$ equals the minimum length of the elements $w\in W$ such that $w(x_{0})=x$ by [19, Corollary 2.13].
We may associate a quantum Hafnian and a Poincaré polynomial to a general quasiparabolic set by defining $\text{\rm QHf}(X)=\sum_{x\in X}q^{\lambda(x)}x$ and $PS_{X}(q)=\sum_{x\in X}q^{\lambda(x)}.$ However, we will not require this level of generality, because our focus for the rest of the paper will be on quasiparabolic sets that arise from sets of orthogonal roots via the following construction.
@@ 3.2. Motivating examples
We now present the two motivating examples of Definition 3.2: one based on perfect matchings, and the other based on permutations. The following definition will be helpful in the context of the first example.
@@
Let $M=\{\{i_{1},j_{1}\},\ldots,\{i_{k},j_{k}\}\}$ be a perfect matching of the set $[2k]$, where we have $k\geq 2$, $i_{1}<\ldots<i_{k}$, and $i_{t}<j_{t}$ for all $1\leq t\leq k$. Following [17, §2], we define $\sigma_{M}\in S_{2k}$ to be the permutation whose two-line notation is given by
@@
We define $\tau_{M}\in S_{2k}$ to be the fixed-point-free involution whose disjoint cycle notation is given by
@@
    $\{(i,j):j<i\text{\ and\ }u_{ji}\in\mathscr{D}_{U}(R(M))\}=\{(i,j):(j,i)\in D_{M}\}$
    is the Rothe diagram $\hat{D}_{\tt FPF}(\tau_{M})$ of the fixed-point-free involution $\tau_{M}$ in the sense of Hamaker–Marberg–Pawlowski [13, (3.3)].
@@
    The generalized quantum Hafnian $\text{\rm QHf}(U)$ is the $q$-Hafnian in the sense of Jing–Zhang [18, Proposition 3.6], and the generalized quantum Pfaffian $\text{\rm QPf}(U)$ is the $q$-Pfaffian in the sense of Strickland [21] and Jing–Zhang [18, Proposition 3.4].
@@
    The generalized Hafnian $\text{\rm Hf}(U)$ is the Hafnian of a $2k\times 2k$ symmetric matrix $(a_{ij})$ satisfying $a_{ij}=u_{ij}$ for all $1\leq i<j\leq 2k$, and the generalized Pfaffian $\text{\rm Pf}(U)$ is the Pfaffian of a $2k\times 2k$ skew-symmetric matrix $(a_{ij})$ satisfying $a_{ij}=u_{ij}$ for all $1\leq i<j\leq 2k$.
@@
If $\beta,u_{ij}\in U$ satisfy $s_{\beta}(u_{ij})<0$, then we must either have $\beta=u_{aj}$ for some $a\leq i$, or $\beta=u_{ib}$ for some $b\leq j$. It follows that $u_{ij}$ dominates $R(M)$ if and only if we have both $\tau_{M}(i)>j$ and $\tau_{M}(j)>i$, proving (iii).
@@
is the Rothe diagram $\hat{D}_{\tt FPF}(\tau)$ of the fixed-point-free involution $\tau$ as defined by Hamaker–Marberg–Pawlowski [13, (3.3)]. Part (iv) follows by applying (iii) to this definition.
It follows from [13, Proposition 3.6] that the level, $\rho_{U}(R(M))=|\hat{D}_{\tt FPF}(\tau_{M})|$, of $R$ is equal to $(\ell(\tau_{M})-k)/2$. Lemma 3.10 (iii) implies that $\rho_{U}(R(M))=\ell(\sigma_{M}$), and this completes the proof of (v).
Rains–Vazirani [19, §4] proved that the function $(\ell(\tau_{M})-k)/2$ is a quasiparabolic level function for the set of fixed-point-free involutions of $S_{2k}$, regarded as an $S_{2k}$-set in the natural way. Combining this result with part (v) proves (i).
@@
where the sum is over all perfect matchings of $[2k]$. This agrees with Jing–Zhang’s construction of the quantum Hafnian from [18, Proposition 3.6] All the assertions of (vi) and (vii) now follow from the relevant definitions.
Part (viii) is a standard result that has been rediscovered several times; see [12, Proposition 7.1]. ∎
@@
In certain cases, the product of the elements in $\mathscr{D}_{U}(R(M))$ associated to a perfect matching $M$, when regarded as a polynomial in the $\varepsilon_{i}$, has an interpretation in terms of Schubert polynomials; see [13, Theorem 1.3].
@@
    The generalized quantum Hafnian $\text{\rm QHf}(U)$ is the $q$-permanent of the matrix $A=(v_{ij})$ in the sense of [17, (2.7)], and the generalized quantum Pfaffian $\text{\rm QHf}(U)$ is the quantum row permanent $\text{\rm per}_{q}(A)$ of $A$ in the sense of [18, (3.18)].
@@
    The generalized Hafnian $\text{\rm Hf}(U)$ and the generalized Pfaffian $\text{\rm Pf}(U)$ are the permanent $\text{\rm per}(A)$, and the determinant $\det(A)$, respectively.
@@
Parts (ii) and (iii) follow by restriction from Theorem 3.11 (ii) and (iii), with each $k$-root $R(\pi)$ defined in (ii) corresponding to the perfect matching $M\{\{i,\pi(i)+k\}:i\in[k]\}$. Part (iv) then follows from the definition of the Rothe diagram $D(\pi)$ [13, (2.7)]. Part (v), which is well known, follows from the fact that there is a bijection between the inversions of $\pi$ and the Rothe diagram of $\pi$ given by $(i,j)\mapsto(i,\pi(j))$.
@@
By [19, Theorem 3.1], this makes $S_{k}$ into a quasiparabolic set with height function $\ell$, proving (i).
@@
    It is sometimes helpful to identify $U$ as the positions on a $k\times k$ board and $\Omega_{U}$ with configurations of $k$ nonattacking rooks on the board, with each root $v_{ij}=\varepsilon_{i}\pm\varepsilon_{j+k}$ corresponding to the $(i,j)$-th position. A position on the board then dominates a $k$-root if and only if it is vacant and not attacked either from the left or from above. Plotting a $k$-root $R$ and the set $\mathscr{D}_{U}(R)$ on the board yields the original Rothe diagram of the permutation corresponding to the rook configuration; see Figure 1.
@@
    Instead of regarding $\Omega_{U}$ as an $S_{k}\times S_{k}$-set, we may regard it as an $(S_{k},S_{k})$-biset, as in the proof of part (v) of the theorem. From this perspective, $\Omega_{U}$ has a left $S_{k}$-action permuting the rows, and a right $S_{k}$-action permuting the columns. The actions commute with each other, and $\Omega_{U}$ is a quasiparabolic $S_{k}$-set under each action.